# Explainable Visual Question Answering with Multimodal Rationales

### v2

@@ Abstract
Visual Question Answering (VQA) is a challenging task of predicting the answer to a question about the content of an image. Prior works directly evaluate the answering models by simply calculating the accuracy of the predicted answers. However, the inner reasoning behind the prediction is disregarded in such a “black box” system, and we cannot ascertain the trustworthiness of the predictions. Even more concerning, in some cases, the models predict correct answers despite focusing on irrelevant visual regions or textual tokens. To develop an explainable and trustworthy answering system, we propose a novel model termed MRVQA (Multimodal Rationales for VQA), which provides visual and textual rationales to support its predicted answers. To measure the quality of the generated rationales, we introduce a new metric called vtS (visual-textual Similarity) score from both visual and textual perspectives. Considering the extra annotations distinct from standard VQA, MRVQA is trained and evaluated using samples synthesized from some existing datasets. The extensive experiments across four VQA datasets demonstrate that the proposed approach achieves new state-of-the-art results through additional rationale generation, enhancing the trustworthiness of explainable VQA models. The code will be released after publication.
Index Terms:
Multimodal reasoning, Visual Question Answering, Explanations.
I Introduction
@@
Visual Question Answering (VQA) takes as input an image and a natural language query and predicts the corresponding answer based on the understanding of the provided vision-language content. It is more challenging than other fundamental computer vision tasks (e.g., object detection [1]), since it requires fusion and reasoning cross different modalities. With the rapid developments of deep neural networks, recent approaches [2, 3, 4, 5] achieve promising performance on VQA. Normally, the answer accuracy is employed in most VQA systems to measure the correctness of the predictions. However, the users cannot know how reliable the predicted answers are, purely based on the accuracy number. In contrast, explainable VQA (EVQA) tries to link the predicted answer to the given image-question pair by providing explanations in different formats (e.g., image regions of interest, and textual explanations) to support the answering process simultaneously. This explanatory mechanism enables VQA systems to reason for complex multimodal information fusion and scene understanding.
@@
Significant progress has been made in the field of “explainable artificial intelligence” (XAI) [6], with remarkable success in uncovering the inner workings of AI algorithms and providing insights into how they arrive at their decisions. However, there remains a largely unexplored area in developing human-centric explanations for interactive vision-language tasks, such as VQA. Most of the prevailing methods adopt single-track models to improve the explainability: either visual hints (termed “V-Exp”) or textual explanations (termed “T-Exp”), as illustrated in Fig. 1 (b). V-Exp methods [7, 8] predict the answer to the given image-question pair and further extract the regions of interest by analyzing the heat maps or attention maps in the inner neural network layers. However, the provided information is usually coarse and inaccurate while the image regions of interest are restricted to a region level, which is not human-friendly for understanding the decision-making process. T-Exp methods [9, 10, 11] aim to describe the visual content of the image that helps explaining the answer in natural language. The explanation partially reproduces the reasoning process behind answering the question. However, it still lacks a rich comprehension of the vision-language input when we aim to directly and visually identify the specific target (i.e., the zebra in the green bounding box from Fig. 1 (d)). Another drawback arises from the inconvenience of textual explanations in complex scenes (e.g., multiple zebras lying on the ground or the presence of horses), leading to longer and more detailed descriptions of the relevant objects. Based on the above concerns in mind, we observe that the current single-track EVQA methods cannot satisfy the credibility of an interactive vision-language system.
@@
Figure 1: Illustration of the differences between standard VQA, single-track EVQA, compositional EVQA and our multimodal EVQA. Our model answers the question and simultaneously predicts textual and visual rationales to support the answer.
@@
In fact, prior to this paper, a few works have made significant strides towards EVQA with compositional reasoning. However, these works essentially provide various relationships among present objects with external information (e.g., scene graphs [12]), resulting in a lack of visual recognition capabilities(as shown in Fig. 1 (c)). Specifically, the textual explanations are either chosen from multiple choices or generated using functional programs from a predefined candidate pool. The shortcuts within the external information allow the model to directly utilize these relationships (e.g., spatial relation and object category), which can harm the model’s general performance. In addition, the external information itself is expensive to collect. To our knowledge, few previous studies have explored free-form textual explanations alongside precise visual predictions.
@@
Moreover, the assessment of generated explanation quality remains limited to basic text similarity measures. The existing methods [9, 8, 13] for EVQA employ widely-used metrics for text generation evaluation, such as BLEU [14], ROUGE [15], and CIDEr [16]. Normally, these metrics quantitatively compare the generated text to reference texts from various textual aspects (e.g., n-gram overlap, semantic similarity, word order, and structure), while ignoring the role of visual content in cross-modality tasks. However, an ideal evaluation of multimodal explanations should consider both visual and linguistic perspectives. Unfortunately, neither the qualitative evaluation of attention maps nor the correspondence score between detected bounding boxes and referred expressions effectively enhances the explainability and credibility of VQA.
@@
Therefore, in this work, we attempt to address these problems and move towards a multimodal EVQA task. Specifically, we propose a new model called MRVQA (Multimodal Rationale for Visual Question Answering) that not only predicts accurate answers but also delivers both textual and visual rationales. To highlight the differences between existing EVQA methods and our approach, we present an example in Fig. 1. Compared to single-track ones, our approach can simultaneously provide multimodal rationales, offering more reliable and user-friendly explanations. For instance, the textual rationale “One zebra gazes in the opposite direction” and the visual rationale, represented by the green bounding box, complement each other to support the answer prediction “no”. Different from compositional ones, our approach generates free-form textual rationales and precise bounding boxes without relying on external information. From a data perspective, we synthesize a multimodal rationale dataset by augmenting existing datasets to alleviate the annotation effort required by EVQA. From a model perspective, we base our approach on a Transformer architecture [17], projecting features into the latent space of a large language model for textual predictions, and fusing features to extract bounding boxes by leveraging a pre-trained open-vocabulary detector. Additionally, we propose a loss function designed to promote the effective cross-alignment and balance among multitask predictions. From an evaluation perspective, we introduce a new metric, which incorporates both visual and textual views to assess the generated rationales and advance EVQA research.
@@
  • To assess the quality of generated rationales, we introduce a new metric called vtS for evaluating natural language from a visual perspective. We employ both textual similarity and the visual presence of key objects to enhance the comprehensive evaluation of the explainability.
@@
  • We demonstrate the effectiveness of our model with extensive experiments under different settings. The proposed model achieves superior performance over the existing methods on both single-track and multimodal EVQA datasets.
II Related Work
II-A Visual Question Answering
@@
Visual Question Answering (VQA) has been defined as the task of providing an answer to a question based on the context of an image. As a fundamental and challenging task in the vision-language field, VQA requires cross-modality understanding and reasoning. Traditionally, it is considered as a classification problem with a pre-defined answer candidate pool and later expanded to free-form open-ended answering. As this task takes as input both images and questions, VQA models address the problem through diverse approaches in different stages, including feature extraction [2], alignment [18], fusion [19], and prediction [20]. With the rapid development of Transformers [17], both natural language processing and computer vision communities have witnessed remarkable advances in vision-language tasks [21, 22]. For instance, MCAN [23] introduced a deep modular co-attention network by leveraging intra-modal self-attentions and cross-modal guided attentions, which enhanced the feature representation and alignment. Some recent works [24, 25] investigate leveraging the vision-language pre-training from large-scale data to enhance the efficacy of VQA models and enable complex reasoning, pioneering the burgeoning trend in the vision-language community. These methods rely on extensive pre-training data and meticulous fine-tuning for downstream tasks, yet they still infer answers through a “black-box” system, which diminishes the explainability of results. In contrast, we utilize multimodal rationales to support answer predictions and enhance model credibility.
II-B Explanations for VQA
@@
Deep neural networks are typically “black-box” systems because their internal mechanics and decision-making processes are not easily interpretable or understandable. Some pioneering works have emerged that bring explanations to VQA. Visualization is the first idea adopted to display the inherent details of models, aiding users in understanding the predictions. For instance, VQA-HAT [7] collected human attention maps of where humans choose to look so as to answer questions, visually providing coarse explanations for VQA. Textual explanations utilize natural language to articulate the reasoning behind the system’s predictions. VQA-X [8] combined textual explanations annotated by humans and visual pointing (similar to attention maps) to construct a multimodal reasoning dataset. Subsequently, Li et al. [9] further emphasized textual explanations by employing a semi-automatic construction method using caption annotations, creating a large-scale EVQA dataset (i.e., VQA-E). VCR [26] transferred the generation task to a simpler classification one by building a multiple choice movie-scene dataset. GQA-REX [13] constructed a large-scale visually grounded textual explanation dataset by extending the previous GQA [27] dataset with an automatic scheme. However, the displayed decision-making process heavily relies on the pre-defined functional programs with external information (e.g., scene graphs and correlations), which reduces its applicability in a diverse open world. Chen et al. introduced a visually grounding dataset VizWiz-VQA-Grounding [28] with questions asked by people with visual impairments. Although the queried objects were annotated with boundaries, the questions lack both in quantity and complexity, and there are no accompanying explanations for the corresponding answers. Differently, we synthesize an EVQA dataset by augmenting existing datasets and provide multimodal explanations through free-form texts and precise bounding boxes.
II-C Generating Multimodal Explanations
@@
The increasing interest in understanding the reasoning behind VQA has led to the development of various methods across both textual and visual perspectives [26, 13]. Park et al. [8] first adopted an LSTM [29] to generate textual explanations and subsequently displayed the internal visual focus through attention maps. Li et al. [9] employed two prediction heads to generate answers and textual explanations, respectively. Wu et al. [10] proposed to leverage the consistency between gradient-based visual explanations and answers to generate robust textual explanations. These earlier works employ encoders to separately extract intra-modality features, which are then simply fused being passed to the final prediction stage. Marasović et al. [30] incorporated pretrained language models with object recognition, visual semantics, and scene graphs to generate textual explanations. Dua et al. [3] introduced a network comprising a generation module and a refinement module, both utilizing sequential LSTMs for answer and explanation generation. DMRFNet [11] leveraged pre-trained semantic relation embeddings and adaptively fused spatial and semantic features through multi-graph reasoning and fusion layers. REX [13] enhanced cross-modality alignment by considering the semantic similarity between words and image regions. Subsequently, VCIN [31] extended relationships to include words, image regions, and explanation tokens through a gating transformer with variational causal inference. However, these methods restrict the relations with pre-defined compositions (e.g., semantic relations, scene graphs and functional programs). In addition, visual explanations are usually generated through gradient-based segmentation proposals [10] or post-processed visual regions [13, 31]. Different from the aforementioned methods, our proposed approach focuses on generating free-form textual explanations and object-level visual explanations.
II-D Grounding with Text
Referring expression comprehension, also known as visual grounding [32, 33], is a task of locating the targets conditioned on the provided natural language expression. Existing models can be categorized into two groups: two-stage methods [34] first generate region proposals and then select the best matching results while one-stage methods [35] directly predict the detected results. For instance, Grounding-DINO [1] concatenated all category names into a single input text, encoded to a text embedding, and then output the top-k highest scores for object detection. However, these visual grounding models rely on specific and direct text inputs (e.g., category) to identify target objects, whereas the input questions may not always provide this information. For example, when posed with the question “which one is taller in the park?”, the aforementioned methods might struggle to accurately extract the relevant objects. In this work, we address this problem by generating textual rationales that explain the predictions by detailing the involved targets. It is worth mentioning that we have conducted an ablation in Section IV-C to analyze the impact of linguistic inputs.
III Method
III-A Problem Formulation
We first revisit the standard VQA setting. Given an image $I$ and a natural language question $Q$, the desired output of a VQA model is an answer $A$ (i.e., a single word or a short phrase) that relates the queried image. As reviewed in Section II-A, standard VQA methods provide direct answers but lack explainability and credibility.
@@ III-A Problem Formulation
To overcome the limitation in standard VQA, we expand the model’s predictions beyond a single answer to include multimodal rationales, offering explanations from both visual and textual perspectives. Regarding the textual part, we follow previous single-track EVQA methods [9, 11] and employ a similar text description to explain the related context for answering the question, which we refer to as the textual rationale $TR$. In contrast to previous methods [36, 8] that use region-level attention maps for visual explanations, we utilize precise bounding boxes of relevant objects as the visual rationale $VR$. This setting directly links the question and answer through clear and object-specific visual information. To this end, given $I$ and $Q$, the goal of our multimodal EVQA model is to predict multiple results: (1) an answer $A$; (2) a textual rationale $TR$; and (3) a visual rationale $VR$. It is simply formulated as follows,
@@
where $\theta$ represents the parameters of an EVQA model. With the support of both visual and textual rationales, users can easily understand the logical context behind the answering and see the visual correspondence, as displayed in Fig. 1 (d). It is particularly valuable for users who wish to explore more details in an interactive scenario. Moreover, these rationales enhance the model’s explainability and credibility.
III-B Multimodal EVQA Dataset Synthesis
@@
To perform the multimodal EVQA task, we review the available EVQA datasets [7, 9, 8, 26, 13] as discussed in Section II-B. Unfortunately, we have observed that these datasets lack either precise visual evidences or related textual descriptions. To avoid the labor-intensive process of annotating explanation-aware VQA examples, we propose a semi-automatic approach to synthesize examples by augmenting an existing EVQA dataset. We select the VQA-E [9] dataset for its comprehensive set of examples that include detailed textual explanations and, importantly, for its integration with the COCO dataset [37] that provides precise bounding-box annotations for object detection. Briefly, we complement the visual explanations by incorporating these annotations. We begin by aligning potential annotations with samples from the VQA-E dataset based on shared objects. Subsequently, we manually refine these results to ensure the dataset quality.
Figure 2: The overall process of the multimodal EVQA dataset synthesis. The text in the brackets shows an example.
@@
Figure 2 illustrates the overall process involved in synthesizing the multimodal EVQA dataset. Specifically, we first extract all nouns from the question-answer-text triplets in the VQA-E dataset. These nouns usually identify the objects relevant to the explanation of answer predictions. For instance, given a question “What color is the tie of that boy?”, the visual focus of the explanation corresponds to the objects “tie” and “boy”. However, analyzing all nouns is impractical due to the presence of over 10K categories. To address this, we conduct a frequency analysis of the nouns, selecting those that appear more than 20 times. This step yields 1,247 distinct nouns that can be matched with the visual labels. Unfortunately, COCO only includes annotations for 80 object categories. To address this, we implement a strategy that classifies candidate nouns into broader categories aligned with the existing COCO annotations. For instance, nouns like “man”, “woman”, “people”, “boy”, and “girl” are categorized under the “Person” category in COCO. However, not all bounding boxes within a given category are relevant for explanations of a specific image-question pair. For example, in Fig. 2, only one sheep is applicable as the visual rationale for the question about the “right sheep”, despite two sheep present. Additionally, the extracted objects from nouns may not fully capture the logical context of the answering process. For instance, with a question “What is the man doing?” where the answer is “surfing” and the text is “a man on the wave” we can only extract the bounding box for the noun “man”. Ideally, the visual rationale should encompass both “man” and “surfboard”. To ensure the accuracy and relevance of visual rationales, we manually refine the bounding boxes by removing irrelevant objects and incorporating additional objects that are essential for the explanations. This process results in the synthesis of a dataset, including image-question pairs annotated with answers, textual rationales, and visual rationales, thereby facilitating multimodal EVQA.
TABLE I: Statistics for the comparison between the synthesized EVQA dataset and prior datasets. Bounding box and scene graph are abbreviated as B-box and S-G, respectively.
Dataset Image #Image #Q&A #TR #VR VR Format
VQA-HAT [7] COCO 20K 59K - 62K Region
VQA-X [8] COCO 28K 32K 41K 6K Region
VQA-E [9] COCO 108K 269K 269K - -
VCR [26] Movies 99K 264K 264K - -
Vizwiz-G [28] Phone 9K 9K - 9K Boundary
GQA-REX [13] GQA 82K 1040K 1040K 82K S-G
VQA-E-Syn COCO 20,367 33,726 33,726 93,377 B-box
We further analyze the synthesized multimodal EVQA dataset (VQA-E-Syn) and compare it with available datasets in Table I. Overall, the dataset comprises 33,726 question-answer pairs derived from 20,367 images, adhering to the original split in the VQA-E dataset. Each pair is accompanied by a textual rationale and an average of 2.8 bounding boxes for visual rationales. The distinctions from previous EVQA datasets are highlighted in the last three columns (i.e., the quantity and format of multimodal rationales). For instance, in comparison to the VQA-X dataset [8], our VQA-E-Syn dataset provides more comprehensive and precise visual rationales. We believe this synthesized dataset will foster future research in multimodal EVQA, which will be made publicly available.
@@
III-C Multimodal EVQA Network
The overall framework of the proposed MRVQA model is shown in Fig. 3. Since our intention is to establish a simple and computation-friendly baseline model for the multimodal EVQA task, we have designed a model to be clean and adaptable to various scenarios. Here, we introduce the core components of our approach: input representation, encoder-decoder architecture, projection module, predictors, and loss functions.
III-C1 Input Representation
Different from previous methods that use separate visual and linguistic feature extractors, we utilize a pre-trained vision-language model (i.e., CLIP [38]) to simultaneously represent both the input image and question. The CLIP model, pre-trained on large-scale image-caption pairs and benefited from transferable representations, outperforms intra-encoders trained solely on in-domain data. To support the cross-modal understanding required for VQA, we employ the pre-trained CLIP model that leverages powerful encoders (e.g., ViT [39] and Transformers [17]), which enhance robust feature extraction and offer promising scalability.
III-C2 Encoder-Decoder
To effectively understand the queried focus from the visual and linguistic features, it is essential to exploit cross-modality fusion for VQA tasks. To enhance representations without introducing a new backbone or altering the model’s structure, we build on the popular VQA baseline MCAN [23], which cascades several Modular Co-Attention (MCA) layers. The MCA layer comprises self-attention units and guided-attention units. The former capture intra-modal features, while the latter enable cross-modal interactions. We employ the default structure and parameters in MCAN-base to build a simple and clean model for our EVQA task.
III-C3 Projection Module
Large language models (LLMs) have been employed to generate sequential texts in our approach. However, there is a gap between pre-trained representations and LLMs in multimodal tasks. To alleviate the mismatch between different feature spaces, a projection module is often employed to facilitate better adaptation and integration. A straightforward linear projection is widely adopted and has proven effective. To strengthen the connection between pre-trained CLIP features and LLMs, we propose a Transformer-based [17] projection module. The self-attention mechanism enables the learning of global dependencies between tokens, even within the context of long sentences. Inspired by Darcet et al. [40], we introduce two additional tokens as learnable constants for the CLIP features $\underline{I}$ and $\underline{Q}$, respectively. These constant tokens act as registers, allowing the model to store and retrieve meaningful information efficiently. For the entire projection module, we utilize eight self-attention layers to process the dual-stream visual and linguistic features. The projected features $I_{proj}$ and $Q_{proj}$ are calculated as follows,

### v3

# Multimodal Rationales for Explainable Visual Question Answering

@@ Abstract
Visual Question Answering (VQA) is a challenging task of predicting the answer to a question about the content of an image. Prior works directly evaluate the answering models by simply calculating the accuracy of predicted answers. However, the inner reasoning behind the predictions is disregarded in such a “black box” system, and we cannot ascertain the trustworthiness of the predictions. Even more concerning, in some cases, these models predict correct answers despite focusing on irrelevant visual regions or textual tokens. To develop an explainable and trustworthy answering system, we propose a novel model termed MRVQA (Multimodal Rationales for VQA), which provides visual and textual rationales to support its predicted answers. To measure the quality of generated rationales, a new metric vtS (visual-textual Similarity) score is introduced from both visual and textual perspectives. Considering the extra annotations distinct from standard VQA, MRVQA is trained and evaluated using samples synthesized from some existing datasets. Extensive experiments across three EVQA datasets demonstrate that MRVQA achieves new state-of-the-art results through additional rationale generation, enhancing the trustworthiness of the explainable VQA model. The code and the synthesized dataset are released under https://github.com/lik1996/MRVQA2025.
1 Introduction
@@
Visual Question Answering (VQA) takes as input an image and a natural language query and predicts the corresponding answer based on the understanding of the provided vision-language content. It is more challenging than other fundamental computer vision tasks (e.g., object detection [29]), since it requires fusion and reasoning cross different modalities. With the rapid developments of deep neural networks, recent approaches [2, 13, 20, 31] achieve promising performance on VQA. Normally, the answer accuracy [3] is employed in most VQA systems to measure the correctness of the predictions. However, users cannot know how reliable the predicted answers are, purely based on the accuracy numbers. In contrast, explainable VQA (EVQA) tries to link the predicted answer to the given image-question pair by providing explanations in different formats (e.g., image regions of interest and textual explanations) to support the answering process. This explanatory mechanism enables VQA systems to reason for multimodal understanding.
@@
Significant progress has been made in the field of “explainable artificial intelligence” (XAI) [40], uncovering the inner workings of AI algorithms and providing insights into how they arrive at their decisions. However, there remains a largely unexplored area in developing human-centric explanations for interactive vision-language tasks, such as VQA. Most of prevailing methods adopt single-track models to improve the explainability: either visual hints (termed “V-Exp”) or textual explanations (termed “T-Exp”), as illustrated in Fig. 1 (b). V-Exp methods [11, 34] predict answers and further extract the regions of interest by analyzing the heat maps or attention maps in the inner neural network layers. However, the provided information is usually coarse and inaccurate while they are restricted to a region level, which is not human-friendly for understanding the decision-making process. T-Exp methods [23, 46, 50] aim to describe the visual content of the image that helps explaining the answer in natural language. The explanation partially reproduces the reasoning process behind answering. However, it still lacks a rich comprehension of vision-language inputs when we aim to directly and visually identify the specific target (i.e., the zebra in the green bounding box from Fig. 1 (d)). Another drawback arises from the inconvenience of textual explanations in complex scenes (e.g., multiple zebras lying on the ground or the presence of horses), leading to longer and more detailed descriptions of the relevant objects. Based on the above concerns in mind, we observe that the current single-track EVQA methods cannot satisfy the credibility of an interactive vision-language system.
@@
Figure 1: Illustration of the differences between standard VQA, single-track EVQA, compositional EVQA and our multimodal EVQA. Our MRVQA answers the question and simultaneously predicts textual and visual rationales to support the answer.
@@
In fact, prior to this paper, a few works [8, 47] have made significant strides towards EVQA with compositional reasoning. However, these works essentially rely on various relationships among objects with external information (e.g., scene graphs [36]), resulting in a lack of visual recognition capabilities (as shown in Fig. 1 (c)). Specifically, the textual explanations are either chosen from multiple choices or generated using functional programs from a predefined candidate pool. The shortcuts within the external information allow the model to directly utilize these relationships (e.g., spatial relation and object category), which can harm the model’s general performance. In addition, the external information itself is expensive to collect. To our knowledge, few previous studies have explored free-form textual explanations alongside precise visual predictions.
@@
Moreover, the assessment of generated explanation quality remains limited to basic text similarity measures. The existing methods [23, 34, 8] for EVQA employ widely-used metrics for text generation evaluation, such as BLEU [33], ROUGE [26], and CIDEr [44]. Normally, these metrics quantitatively compare the generated text to reference texts from various textual aspects (e.g., n-gram overlap, semantic similarity, word order, and structure), while ignoring the role of visual content in cross-modality tasks. However, an ideal evaluation of multimodal explanations should consider both visual and linguistic perspectives.
@@
Therefore, in this work, we attempt to address these problems with a multimodal EVQA model. Specifically, we propose a new model called MRVQA (Multimodal Rationale for Visual Question Answering) that not only predicts accurate answers but also delivers both textual and visual rationales. To highlight the differences between existing EVQA methods and our approach, an example is presented in Fig. 1. Compared to single-track ones, our approach can simultaneously provide multimodal rationales, offering more reliable and user-friendly explanations. For instance, the textual rationale “One zebra gazes in the opposite direction” and the visual rationale, represented by the green bounding box, complement each other to support the answer prediction “no”. Different from compositional ones, our approach generates free-form textual rationales and precise bounding boxes without relying on external information. From a data perspective, we synthesize a multimodal rationale dataset by augmenting existing datasets to alleviate the annotation effort required by EVQA. From a model perspective, we base our approach on a Transformer architecture [43], projecting features into the latent space of a large language model for textual predictions, and fusing features to extract bounding boxes by leveraging a pre-trained open-vocabulary detector. Additionally, we propose a loss function designed to promote the effective cross-alignment and balance among multitask predictions. From an evaluation perspective, a new metric is introduced, which incorporates both visual and textual views to assess generated rationales and advance EVQA research.
@@
  • To assess the quality of generated rationales, we introduce a new metric called vtS for evaluating natural language from a visual perspective.
@@
  • We demonstrate the effectiveness of our model with extensive experiments under different settings. The proposed model achieves superior performance over existing methods on three EVQA datasets.
2 Related Work
2.1 Visual Question Answering
@@
Visual Question Answering (VQA) has been defined as the task of providing an answer to a question about an image. As this task takes as input both images and questions, VQA models address the problem through diverse approaches in different stages, including feature extraction [2], alignment [28], fusion [5], and prediction [17]. With the rapid development of Transformers [43], both natural language processing and computer vision communities have witnessed remarkable advances in vision-language tasks [37, 22]. For instance, MCAN [48] introduced a deep modular co-attention network by leveraging intra-modal self-attentions and cross-modal guided attentions, which enhanced the feature representation and alignment. Some recent works [30, 51] investigate leveraging the vision-language pre-training from large-scale data to enhance the efficacy of VQA models and enable complex reasoning, pioneering the burgeoning trend in the vision-language community. These methods rely on extensive pre-training data and meticulous fine-tuning for downstream tasks, yet they still infer answers through a “black-box” system, which diminishes the explainability of results.
2.2 Explanations for VQA
@@
Deep neural networks are typically “black-box” systems because their internal mechanics and decision-making processes are not easily interpretable or understandable. Some pioneering works have emerged that bring explanations to VQA. Visualization is the first idea adopted to display the inherent details of models, aiding users in understanding the predictions. For instance, VQA-HAT [11] collected human attention maps of where humans choose to look so as to answer questions, visually providing coarse explanations for VQA. Textual explanations utilize natural language to articulate the reasoning behind the system’s predictions. VQA-X [34] combined textual explanations annotated by humans and visual pointing (similar to attention maps) to construct a multimodal reasoning dataset. Subsequently, Li et al. [23] further emphasized textual explanations by employing a semi-automatic construction method using caption annotations, creating a large-scale EVQA dataset (i.e., VQA-E). VCR [49] transferred the generation task to a simpler classification one by building a multiple choice movie-scene dataset. GQA-REX [8] constructed a large-scale visually grounded textual explanation dataset by extending the previous GQA [16] dataset with an automatic scheme. However, the displayed decision-making process heavily relies on the pre-defined functional programs with external information (e.g., scene graphs and correlations), which reduces its applicability in a diverse open world.
2.3 Generating Multimodal Explanations
@@
The increasing interest in understanding the reasoning behind VQA has led to the development of various methods across both textual and visual perspectives [49, 8]. Park et al. [34] first adopted an LSTM [15] to generate textual explanations and subsequently displayed the internal visual focus through attention maps. Li et al. [23] employed two prediction heads to generate answers and textual explanations, respectively. Wu et al. [46] proposed to leverage the consistency between gradient-based visual explanations and answers to generate robust textual explanations. These earlier works employ encoders to separately extract intra-modality features, which are then simply fused being passed to the final prediction stage. Marasović et al. [32] incorporated pretrained language models with object recognition, visual semantics, and scene graphs to generate textual explanations. DMRFNet [50] leveraged pre-trained semantic relation embeddings and adaptively fused spatial and semantic features through multi-graph reasoning and fusion layers. REX [8] enhanced cross-modality alignment by considering the semantic similarity between words and image regions. Subsequently, VCIN [47] extended relationships to include words, image regions, and explanation tokens through a gating transformer with variational causal inference. However, these methods restrict the relations with pre-defined compositions (e.g., semantic relations, scene graphs and functional programs). In addition, visual explanations are usually generated through gradient-based segmentation proposals [46] or post-processed visual regions [8, 47]. Different from the aforementioned methods, our proposed approach focuses on generating free-form textual explanations and object-level visual explanations.
3 Method
3.1 Problem Formulation
@@
We first revisit the standard VQA setting. Given an image $I$ and a natural language question $Q$, the desired output of a VQA model is an answer $A$ (i.e., a single word or a short phrase) that relates the queried image. To overcome the limitation in standard VQA, we expand the model’s predictions beyond a single answer to include multimodal rationales, offering explanations from both visual and textual perspectives. Regarding the textual part, we follow previous single-track EVQA methods [23, 50] and employ a similar text description to explain the related context for answering the question, which we refer to as the textual rationale $TR$. In contrast to previous methods [7, 34] that use region-level attention maps for visual explanations, we utilize precise bounding boxes of relevant objects as the visual rationale $VR$. To this end, given $I$ and $Q$, the goal of our multimodal EVQA model is to predict multiple results: (1) an answer $A$; (2) a textual rationale $TR$; and (3) a visual rationale $VR$. It is simply formulated as follows,
@@
where $\theta$ represents the parameters of an EVQA model. It is particularly valuable for users who wish to explore more details in an interactive scenario. Moreover, these rationales enhance the model’s explainability and credibility.
3.2 Multimodal EVQA Dataset Synthesis
@@
Unfortunately, we have observed that the available EVQA datasets [11, 23, 34, 49, 8] lack either precise visual evidences or related textual descriptions. To avoid the labor-intensive process of annotating explanation-aware VQA examples, we propose a semi-automatic approach to synthesize examples by augmenting an existing EVQA dataset. We select VQA-E [23] for its comprehensive set of examples that include detailed textual explanations and, importantly, for its integration with COCO [27] that provides precise bounding-box annotations for object detection.
@@
Specifically, we first extract all nouns from the question-answer-text triplets in the VQA-E dataset. These nouns usually identify the objects relevant to the explanation of answer predictions. However, analyzing all nouns is impractical due to the presence of over 10K categories. To address this, we conduct a frequency analysis of the nouns, selecting those that appear more than 20 times. This step yields 1,247 distinct nouns that can be matched with the visual labels. Unfortunately, COCO only includes annotations for 80 object categories. To address this, we implement a strategy that classifies candidate nouns into broader categories aligned with the existing COCO annotations. For instance, nouns like “man”, “woman”, “people”, and “girl” are categorized under the “Person” category in COCO. However, not all bounding boxes within a given category are relevant for explanations of a specific image-question pair. For example, although two sheep are present in an image, only one serves as the correct visual rationale for a question concerning the “right sheep”. Additionally, noun-based object extraction may miss key contextual elements. For example, in answering “What is the man doing?” with ground truth “surfing” and “a man on the wave”, only “man” yields a bounding box, while the full rationale should include both the man and the surfboard. To ensure the accuracy and relevance of visual rationales, two annotators, following detailed guidelines, manually refined the bounding boxes over two weeks by removing irrelevant objects and adding those essential to the explanation. This process results in the synthesis of a dataset (VQA-E-Syn). Overall, it comprises 33,726 question-answer pairs derived from 20,367 images, adhering to the original split in VQA-E. Each pair is accompanied by a textual rationale and an average of 2.8 bounding boxes for visual rationales.
@@
3.3 Multimodal EVQA Network
The overall framework of the proposed MRVQA model is shown in Fig. 2. Here, the core components of our approach are introduced as follows.
Input Representation. Different from previous methods that use separate visual and linguistic feature extractors, we utilize a pre-trained vision-language model (i.e., CLIP [39]) to simultaneously represent both the input image and question. The CLIP model, pre-trained on large-scale image-caption pairs and benefited from transferable representations, outperforms intra-encoders trained solely on in-domain data. With the pre-trained CLIP model, we generate robust visual and linguistic features for our EVQA task.
Encoder-Decoder. To effectively understand the queried focus from the visual and linguistic features, it is essential to exploit cross-modality fusion for VQA tasks. To enhance representations without introducing a new backbone or altering the model’s structure, we build on the popular VQA baseline MCAN [48], which cascades several Modular Co-Attention (MCA) layers. The MCA layer comprises self-attention units for intra-modal interaction and guided-attention units for cross-modal interaction.
Projection Module. To alleviate the mismatch between the pre-trained CLIP and our textual rationale generator (Large Language Models, LLMs described in the following part), we employ a Transformer-based [43] projection module, facilitating multimodal adaptation and integration. The self-attention mechanism enables the learning of global dependencies between tokens, even within the context of long sentences. Inspired by Darcet et al. [10], we introduce two additional tokens as learnable constants for the CLIP features $\underline{I}$ and $\underline{Q}$, respectively. These constant tokens act as registers, allowing the model to store and retrieve meaningful information efficiently. For the entire projection module, we utilize eight self-attention layers to process the dual-stream visual and linguistic features. The projected features $I_{proj}$ and $Q_{proj}$ are calculated as follows,